The arithmetic average growth rate in sales is 7%, the geometric average growth rate in sales is 6% and the regression equation to estimate sales is Sales = 50 + 15(T). In 2017 (T=19), sales were 335.

1. What is the forecasted sales for 2018 using the geometric average? (express your answer to 2 decimal places)

2. What is the forecasted sales for 2018 using the arithmetic average? (express your answer to 2 decimal places)

3. What is the forecasted sales for 2018 using the regression equation?
Answer #1

Sales in 2017 = 335

Arithmetic average growth rate in sales = 7%

Geometric average growth rate in sales = 6%

1.

Forecasted sales for 2018 using geometric average = (1+0.06)*335

Forecasted sales for 2018 using geometric average = 355.1

2.

Forecasted sales for 2018 using arithmetic average = (1+0.07)*335

Forecasted sales for 2018 using arithmetic average = 358.45

3.

Forecasted sales for 2018 using the regression equation

Sales = 50 + 15(T)

For 2017, T = 19

For 2018, T = 20

Sales = 50 + 15*20

Sales = 350